OpenStudy (anonymous):

3sqrt{x^-6} 3sqrt{x^-6} @Mathematics

OpenStudy (anonymous):

\[\sqrt[3]x ^{-6}{?}\]

OpenStudy (alfie):

\[3\sqrt {{x^{( - 6)}}} \to 3\sqrt {\frac{1}{{{x^6}}}} \to 3\sqrt {\frac{{{1^6}}}{{{x^6}}}} \to \frac{3}{{{x^3}}}\]

OpenStudy (anonymous):

\[3\sqrt{x^{-6}} = 3x^{-3} = \frac{3}{x^3}\]

OpenStudy (anonymous):

would any of these be the answers that similar to your answer -(x^2) or (1/x^0.500) or (1/x^2) or (x^2) i had chosen the first one. is that right?

OpenStudy (anonymous):

\[\frac{1}{x^{0.500}} = \frac{1}{x^{\frac{500}{1000}}} = \frac{1}{x^{\frac{1}{2}}} = \frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

oh so I was wrong? its the second on.e

OpenStudy (anonymous):

3/x^3

OpenStudy (anonymous):

hmm thats not one of the answers they gave me to choose from i dont know then